Could you please help me on these last 5 questions that I can't seem to figure out??
1. Find the maximum or minimum of the following quadratic function: y =x^2 - 10x + 25.
2. Find the maximum or minimum of the following quadratic function: y =x^2 - 2x - 48.
3. Find the maximum or minimum of the following quadratic function: y = x^2 - x - 56.
4. Find the maximum or minimum of the following quadratic function: y = 16x^2+ 40x + 25.
5. Find the maximum or minimum of the following quadratic function: y = x^2 - 16.

Answer by lwsshak3(11628)

1. Find the maximum or minimum of the following quadratic function: y =x^2 - 10x + 25.
2. Find the maximum or minimum of the following quadratic function: y =x^2 - 2x - 48.
3. Find the maximum or minimum of the following quadratic function: y = x^2 - x - 56.
4. Find the maximum or minimum of the following quadratic function: y = 16x^2+ 40x + 25.
5. Find the maximum or minimum of the following quadratic function: y = x^2 - 16.
**
All the problems above require the same method (completing the square) to solve.
Standard form of equation for the problems: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex.
If the lead coefficient is >0, parabola open up, and has a minimum. If lead coefficient <0, parabola open down, and has a maximum. The x-term of the vertex is the maximum or minimum value.
I will do the first 2, and let you do the rest.
..
1) y =x^2 - 10x + 25
completing the square
y =(x^2-10x+25)-25 + 25
y=(x-5)^2+0
vertex: (5,0)
minimum=0
..
2) y =x^2 - 2x - 48
completing the square
y =(x^2-2x+1)-1-48
y=(x-1)^2-49
vertex: (1,-49)
minimum=-49